the repulsive force between the north poles of 2 magnets is inversely proportional to the square of the distance between the 2 poles. if the repulsive force is 40 pounds when the distance between the 2 poles is 2 inches, what is the repulsive force when the distance between the 2 poles is 4 inches?
Answer by stanbon(75887)

the repulsive force between the north poles of 2 magnets is inversely proportional to the square of the distance between the 2 poles.
-----
f = k/d^2
-----
Solve for "k" using "if the repulsive force is 40 pounds when the distance between the 2 poles is 2 inches"
40 = k/2^2
k = 160
-----
f = 160/d^2
------
What is the repulsive force when the distance between the 2 poles is 4 inches?
--
f(4) = 160/4^2
f(4) = 10 pounds
========================
Cheers,
Stan H.
=====================